204800=50*2^3/D
Note that 3/D is the exponent

Answer by scott8148(6628)

dividing by 50 __ 4096=2^(3/D)

2^12=2^(3/D) __ 12=3/D __ D=3/12


note: 4096 is 4k in computer talk (binary) __ 1k=1024=2^10